OpenStudy (anonymous):

HELP 5 QUESTIONS WILL MEDAL!!!

OpenStudy (anonymous):

Erica purchased adult and youth tickets for a movie. She bought x adult tickets for $7 each and y youth tickets for $5. Write an expression that can be used to show the total cost, C, of all tickets. C = 7x + 5y C = (7 + 5)x C = (7 + 5)(x + y) C = 7y + 5x

OpenStudy (anonymous):

Write an expression that is equivalent to two over three (4x + 9). eight over threex + 6 eight over threex + 9 two over threex + 6 two over threex + 9

OpenStudy (anonymous):

Jason purchased a pack of game cards that was on sale for 13% off. The sales tax in his county is 8%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards. y - 0.13y 1.08(0.87y) 0.08(0.87y) y - 0.87y + 0.08y

OpenStudy (anonymous):

Which expression is equivalent to one over fourn - 16? one over four(n - 4) one over four(n - 64) 4(n - 4) 4(n - 64)

OpenStudy (anonymous):

Write an expression for the perimeter of the triangle shown below: A triangle is shown with side lengths labeled 2.3x plus 14, 2x, and negative 0.2x plus 15. 4.1x + 29 4.1x - 29 4.5x + 29 4.5x - 29

OpenStudy (anonymous):

Erica purchased adult and youth tickets for a movie. She bought x adult tickets for $7 each and y youth tickets for $5. Write an expression that can be used to show the total cost, C, of all tickets. total cost for adult (A) = 7x total cost for youth (Y) = 5y Total cost (C) = A+Y

OpenStudy (anonymous):

an expression that is equivalent to two over three (4x + 9). = 2/3(4x+9) eight over three(x + 6) = 8/3(x+6) eight over threex + 9 = 8/3(x+9) two over threex + 6 = 2/3(x+6) two over threex + 9 = 2/3(x+9) which do you think?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

And whats your reason?

OpenStudy (anonymous):

A triangle is shown with side lengths labeled 2.3x plus 14, 2x, and negative 0.2x plus 15. 2.3x +14 +2x +(-0.2x) +15 so answer is?

OpenStudy (anonymous):

one over fourn - 16 = 1/4(n-16) one over four(n - 4) = 1/4(n-4) one over four(n - 64) = 1/4(n-64) 4(n - 4) 4(n - 64) and answer will be?

OpenStudy (anonymous):

Jason purchased a pack of game cards that was on sale for 13% off. The sales tax in his county is 8%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards. y - 0.13y + 0.08y = final cost

OpenStudy (anonymous):

you are confusing me

OpenStudy (anonymous):

can you just tell me the answers then I will give you a medal!?!

OpenStudy (anonymous):

How about you tell me and then I will give you the answer. Starting from the top

OpenStudy (anonymous):

Tell your answers

OpenStudy (anonymous):

D for 1 and c for 2 also 3 4 and 5 are a b and d

OpenStudy (anonymous):

those are my answers!

OpenStudy (anonymous):

1 = D 2 = C 3 = A 4 = B 5 = D

OpenStudy (anonymous):

@hotdancingbabe14

OpenStudy (anonymous):

@grace111

OpenStudy (anonymous):

@anonymous_user

OpenStudy (anonymous):

1 = A 2 = C 3 = B 4 = D 5 = A

OpenStudy (anonymous):

THANK YOU FINALLY!

OpenStudy (anonymous):

Your welcome